A cross country skier skis 9 km north and 12 km east on a horizontal show field. How far and in what directions is she from the starting point?


1
Expert's answer
2021-09-20T16:32:03-0400


The magnitude of the final displacement can be found from the right triangle in the figure:


"d = \sqrt{d_1^2 + d_2^2} = \sqrt{9^2 + 12^2} =15km"

The direction is the following:


"\theta = \arctan\left( \dfrac{d_1}{d_2}\right)=\arctan\left( \dfrac{9}{12}\right) \approx 37\degree"

Answer. Magnitude: 15 km. Direction: 37 degrees from the East.